# Organizing Unstructured Image Collections using Natural Language

### v1

@@ I.1 Further Details on Discovering and Mitigating Dataset Bias
Implementation Details of Training GroupDRO: To conduct debiased training using GroupDRO (Sagawa et al., 2020), we first used our predicted distribution to define four distinct training groups, rather than relying on the ground-truth distribution. We closely followed the training protocol outlined in B2T (Kim et al., 2024) and GroupDRO (Sagawa et al., 2020). Specifically, we fine-tuned a ResNet-50 (He et al., 2016) model pre-trained on ImageNet (Deng et al., 2009), using the training split of the CelebA dataset (Liu et al., 2015). The training was performed using the SGD optimizer (Ruder, 2016) with a momentum of 0.9, a batch size of 64, and a learning rate of $1\times 10^{-5}$. We applied a weight decay of 0.1 and set the group adjustment parameter to zero. The model was trained over 50 epochs. For evaluation, we reported both the average and worst-group test accuracies, selecting the model from the epoch that achieves the highest worst-group accuracy on the validation set. The final evaluation and comparison results are provided in Fig. 7.
I.2 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
Image Generation for the Subject Occupation: Following prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016), we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator). We used two state-of-the-art T2I diffusion model, DALL·E3 (Betker et al., 2023) and Stable Diffusion (SDXL) (Podell et al., 2024) to generate 100 images for each occupation for our study. This resulted in a total of 1,800 images. For each occupation, we provide some examples of images generated by DALL·E3 in Fig. 20, while provide some examples of images generated by SDXL in Fig. 21. We only used the simple prompt “A portrait photo of a $<$OCCUPATION$>$” for image generation for all occupations and did not include any potential biases in the prompt.
where $\mathcal{D}_{\text{Occupation}}$ represents the generated images for each occupation, $\mathcal{C}^{l}$ denotes the clusters discovered under the $l$-th criterion, and $p(c^{l}|\mathcal{C}^{l})$ is the probability of each cluster under the current distribution. The resulting score is bounded between $\mathcal{H}_{bias}^{l}\in[0,1]$, where 0 indicates no bias towards a specific cluster (concept) under the evaluated criterion, and 1 indicates that the images are completely biased towards a particular cluster (concept) (e.g.,, “Grey” hair color) within the current bias dimension (e.g., Hair color). We used the score defined in Eq. 2 to quantify the biases for each occupation across the 10 discovered grouping criteria. We report the bias intensity score for each occupation and each model across the 10 discovered grouping criteria in Fig. 18.
Human Evaluation Study Details: To assess the alignment between our method’s predictions and human judgments on bias detection, we conducted a user study to gather human evaluation results for the generated images. As shown in the questionnaire example in Fig. 22, participants were presented with images generated by DALL-E3 and SDXL for each occupation and were asked to identify the bias direction (dominant class) for each of the 10 discovered criteria and rate the bias intensity on a scale from 0 to 10. We collected responses from 54 anonymous participants, resulting in 6 human evaluations for each occupation and each criterion.
The Absolute Mean Error (AME) between the bias intensity scores predicted by our system and those rated by humans (scaled to 0 to 1) was 0.1396. Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach. Detailed user study results are provided in Sec. I.2. We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers. In Fig. 18, we present the human evaluation results, averaged across all participants for each model, occupation, and criterion, with the human ratings scaled from 0 to 1.
Complete Results and Additional Findings: In Fig. 18, we present the detailed bias detection results for each model, occupation, and criterion, alongside human evaluation scores for reference. A particularly interesting phenomenon emerges: While DALL·E3 significantly outperforms SDXL on the well-known bias dimensions (e.g.,, Gender, Race, Age, and Skin tone), both DALL·E3 and SDXL exhibit moderate to strong biases along the novel bias dimensions (e.g.,, Hair color, Mood, Attire, and Accessories).
We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” (OpenAI, 2022), designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs. However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied. This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
I.3 Further Details on Analyzing Social Media Image Popularity
Figure 19: Complete analysis of social media photo popularity on the SPID dataset. We display the viral and major popular clusters, along with the popularity distribution of data points within these clusters across all ten discovered criteria (in Grey).
With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems. Traditional research often approaches image popularity as a regression problem (Ortis et al., 2019; Cheng et al., 2024), utilizing metadata like hashtags, titles, or follower counts. However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored. In this study, we applied our proposed method to automatically categorize social media images based on semantic visual elements across different dimensions (criteria). By analyzing these interpretable results alongside image popularity metrics (e.g.,, number of views), we gained insights into the factors contributing to virality and identified common visual traits among popular images. These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
To expand on the discussion in Sec. 5 of the main paper, we present the complete findings across all ten discovered criteria in Fig. 19. Notably, we consistently observed a sharp semantic contrast between the visual elements in viral images and those in the majority of popular images across all ten criteria. For instance, there is a contrast between Urban sophisticated and Modern minimalist under Interior Design, Rustic architecture and Modern architecture under Architecture Style, and Event venues versus Urban residential areas under Location.
This recurring observation reinforces the idea that viral content tends to capture more attention, likely because it features novel, surprising, or striking visual elements. Humans are inherently attracted to stimuli that deviate from the norm (Priester et al., 2004; Bruni et al., 2012; Palmer & Gore, 2014). On the other hand, widely popular yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality. We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
@@ I.3 Further Details on Analyzing Social Media Image Popularity
Appendix J In-depth Discussion of Related Settings
Distinction from Multiple Clustering: Multiple clustering involves finding diverse clusterings from the same dataset in either a semi-supervised (Bae & Bailey, 2006; Ren et al., 2022) or unsupervised (Caruana et al., 2006; Mautz et al., 2020; Yao et al., 2023) manner. However, multiple clustering methods require users to specify the number of clusterings and clusters (or rely on ground-truth annotations for evaluation) (Yu et al., 2024). In other words, $L$ (the number of clusterings) and $K_{l}$ (the number of clusters within each clustering) are assumed as prior knowledge in multiple clustering settings. This strong assumption creates a chicken-or-egg dilemma and is often impractical in real-world applications: Users want to employ multiple clustering methods to understand their data, but how can they predefine the number of valid clusterings and clusters without already having a deep understanding of the entire dataset? Although some strategies exist to determine the number of clusters, most of these methods (Monti et al., 2003; Zhang et al., 2017; Liu et al., 2024e) only work for single clustering. In contrast, SMC requires the model to automatically discover both the number of clusterings and the clusters within them.
Additionally, while multiple clustering methods can reveal diverse sample division patterns and underlying data structures, they do not provide interpretations of the results—specifically, what rules the output clustering follows and the semantic meaning of the clusters. As a result, users often need to manually investigate the clustering outcomes. In stark contrast, SMC methods provide interpretability by describing both the clustering rules and the semantic meanings of the clusters in natural language. This not only offers a more comprehensive understanding of the data but also allows users to combine different clusterings to gain deeper insights into the data distribution.
[…]
Distinction from Multi-label Zero-shot Classification: SMC differs from multi-label zero-shot classification in that the latter (Lee et al., 2018; Huang et al., 2020; Ali & Khan, 2023; Gupta et al., 2023) requires predefined sets of classes under different rules, with the goal being to assign each image to multiple classes from these sets. In contrast, SMC requires discovering the class names (or cluster semantics). In fact, multi-label zero-shot classification can be viewed as a specific instance of SMC when the user explicitly and precisely defines all the clusters and their semantic meanings under different clustering rules.

### v4

@@ Appendix I Failure Case Analysis
Implementation Details of Training GroupDRO: To conduct debiased training using GroupDRO (Sagawa et al., 2020), we first used our predicted distribution to define four distinct training groups, rather than relying on the ground-truth distribution. We closely followed the training protocol outlined in B2T (Kim et al., 2024) and GroupDRO (Sagawa et al., 2020). Specifically, we fine-tuned a ResNet-50 (He et al., 2016) model pre-trained on ImageNet (Deng et al., 2009), using the training split of the CelebA dataset (Liu et al., 2015). The training was performed using the SGD optimizer (Ruder, 2016) with a momentum of 0.9, a batch size of 64, and a learning rate of $1\times 10^{-5}$. We applied a weight decay of 0.1 and set the group adjustment parameter to zero. The model was trained over 50 epochs. For evaluation, we reported both the average and worst-group test accuracies, selecting the model from the epoch that achieves the highest worst-group accuracy on the validation set. The final evaluation and comparison results are provided in Tab. 35.
Appendix Q Why LLMs Improve Image Clustering?
The most compelling aspect of this work lies in our $\mathcal{X}$-Cluster framework’s ability to transform large volumes of unstructured images into natural language and leverage the advanced text understanding and summarization capabilities of LLMs to tackle the challenging Open-ended Semantic Multiple Clustering (OpenSMC) task. This approach draws inspiration from the use of LLMs in the Topic Discovery task within the NLP domain (Eklund and Forsman, 2022). Our core motivation is: “If LLMs can discover topics from documents and organize them, then by converting images into text, we can similarly use LLMs to organize unstructured images.”
Traditional clustering methods (Estivill-Castro, 2002; Caron et al., 2018; Van Gansbeke et al., 2020; Li et al., 2024; Yu et al., 2024) often depend on pre-defined criteria, pre-determined numbers of clusters, fixed feature representations (which require training), and are typically not interpretable. These limitations hinder their applicability to diverse datasets in open-world scenarios, as they demand significant human priors and retraining for each new dataset.
In contrast, LLMs (OpenAI, 2022a, 2023; Touvron et al., 2023; Meta, 2024b, a) excel at understanding, summarizing, and reasoning over high-level semantics expressed in natural language across diverse domains (e.g., everyday content, cultural knowledge, or medical content). Operating in a zero-shot (Kojima et al., 2022), interpretable manner, LLMs are uniquely suited to the SMC task, which aims to discover meaningful and interpretable clustering criteria without requiring prior knowledge or training. By integrating LLMs with MLLMs (Liu et al., 2024b) into the carefully designed $\mathcal{X}$-Cluster framework, we enable the discovery and refinement of clustering criteria directly from the dataset’s content, followed by automatic grouping of the dataset. This design allows our framework to overcome the rigid assumptions of traditional clustering methods, making it automatic, generalizable, and training-free. Our approach provides a novel perspective, demonstrating how clustering tasks can evolve beyond traditional paradigms.
Challenges of employing LLMs to facilitate the SMC task. The main challenge of employing LLMs for the SMC task lies in accurately translating visual content from images into natural language that LLMs can effectively reason with. This is evident from the sensitivity analysis results in App. N: $\mathcal{X}$-Cluster’s performance improves with larger or more powerful MLLMs (see Figure 18 (a)), while it remains relatively insensitive to the specific choice of LLM (see Figure 18 (b)). In other words, the quality of image captions generated by MLLMs is critical for the effective use of LLMs in the SMC task. Specifically, in the first stage of $\mathcal{X}$-Cluster (criteria proposal), captions need to be as comprehensive as possible to provide rich information for LLMs to discover grouping criteria. In the second stage (semantic grouping), criterion-specific captions should precisely capture relevant visual content to provide accurate information for assigning images to clusters.
To enhance caption quality, techniques such as MLLM model ensembling, prompt ensembling (Liu et al., 2024c), or stronger models like GPT-4V (OpenAI, 2023) can improve comprehensiveness. For better precision, advanced prompting methods like CoT (Wei et al., 2022) or FineR (Liu et al., 2024e) can capture nuanced details, while hallucination mitigation tools like Visual Fact Checker (Ge et al., 2024) can reduce noise caused by hallucinations. However, these techniques increase computational costs and framework complexity. In this work, we choose to keep $\mathcal{X}$-Cluster simple yet effective, and we outline these potential improvements for future practitioners.
Appendix R Future Work
Closed-Loop Optimization. In this work, we designed our prompts following the Iterative Prompt Engineering methodology (DeepLearning.AI, 2024) introduced by Isa Fulford and Andrew Ng. In App. E, we provide the exact LLM and MLLM prompts used in our framework and break down each prompt to explain the objectives and purposes behind each design choice. These explanations cover elements such as system prompts, input formatting, task and sub-task instructions, and output instructions. Our focus in this work is on creating a highly generalizable framework, $\mathcal{X}$-Cluster, and we do not perform any closed-loop, dataset-specific prompt optimizations. However, in future work or application scenarios where a labeled training/validation dataset is available, practitioners could build upon our design objectives. By leveraging our proposed evaluation metrics (see § 5) for each step, it would be possible to develop a Open-ended Semantic Multiple Clustering (OpenSMC) system with a closed-loop optimization pipeline to achieve improved performance.
$\bm{\mathcal{X}}$-Cluster on Other Data Types. The core idea of our proposed framework, $\mathcal{X}$-Cluster, is to use text as a proxy (or medium) for reasoning over large volumes of unstructured data, generating human-interpretable insights at scale. As such, $\mathcal{X}$-Cluster can be directly applied to textual data (e.g., documents). Moreover, since natural language is a highly versatile and widely-used medium of representation, $\mathcal{X}$-Cluster can be extended to other data types by converting these data into text (by replacing the captioning module with suitable tools) in future work, such as:
  • Audio Data: Speech-to-Text models like Whisper (Radford et al., 2023) can convert audio data into text, enabling subsequent analysis with $\mathcal{X}$-Cluster.
  • Tabular Data: Table-to-Text models, such as TabT5 (Andrejczuk et al., 2022), can translate tabular data into text, making it compatible with $\mathcal{X}$-Cluster. For tables containing figures, modern MLLMs like LLaVA-Next, which support both OCR and image-to-text capabilities, can handle these elements to create a unified textual representation for $\mathcal{X}$-Cluster.
  • Protein Structures: Protein structure-to-text models, such as ProtChatGPT (Wang et al., 2024a), can convert protein sequences into textual descriptions for analysis with $\mathcal{X}$-Cluster.
  • Point Cloud Data: 3D captioning models, like Cap3D (Luo et al., 2024a), can transform point cloud data or rendered 3D models into text, enabling their analysis using $\mathcal{X}$-Cluster.
We believe the versatile nature of $\mathcal{X}$-Cluster has the potential to open up a broad range of applications across diverse data modalities, fostering new directions in future research.